# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 8.3. Abstract virtually free factors and their Whitehead graphs
Recall that the tree $T$ is bipartite, with all white vertices of degree 2. Thus $T$ is the first barycentric subdivision of certain tree, which we denote $T_{*}$. The vertices of $T_{*}$ are naturally identified with the black vertices of $T$. Our goal is to identify $T_{*}$ with ${\rm Cay}(F,B)$. To do this, we associate to each edge $e$ of $T_{*}$ some orientation and some label from the generating set $B$. Let $t,s$ be the endpoints of $e$, and view them as black vertices of $T$. Let $u$ be the barycenter of $e$, which we view as an appropriate white vertex of $T$ (the unique white vertex of $T$ adjacent to both $t$ and $s$). The constituent spaces at $t$ and $s$ of the tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$, denoted $\Gamma_{t}^{\circ}$ and $\Gamma_{s}^{\circ}$, are the copies of the punctured Whitehead graph $W_{\mathcal{Q}}$. The images in $\Gamma_{t}^{\circ}$ and in $\Gamma_{s}^{\circ}$ of the peripheral space $\Sigma_{u}$ are some $V$-peripherals related to vertices of $\Gamma_{t}$ and $\Gamma_{s}$, and we denote these vertices by $z$ and $y$, respectively. Identifying both $z$ and $y$ as vertices of the Whitehead graph $W_{\mathcal{Q}}$, we have $z,y\in B\cup B^{-1}$, and by the form of the associated $V$-involution $a_{\mathcal{Q}}$, we have $y=z^{-1}$. Without loss of generality, assume that $z\in B$. Orient then the edge $e$ consistently with the order $(z,y)$, and label it with $z$. It is then easy to observe that for any vertex $v$ of $T_{*}$ we have exactly $m$ oriented edges issuing from $v$, labelled bijectively by elements from $B$, and exactly $m$ oriented edges incoming to $v$, also labelled bijectively with elements of $B$. This turns the tree $T_{*}$ to look very similar to the Cayley graph ${\rm Cay}(F,B)$, which is also a tree. To get an explicit identification of $T_{*}$ with ${\rm Cay}(F,B)$, fix a vertex of $T_{*}$ and identify it with the vertex of ${\rm Cay}(F,B)$ corresponding to the unit of $F$. This identification obviously extends to a unique label preserving isomorphism of oriented graphs, and we take this isomorphism as the demanded identification.
Given the identification of $T_{*}$ with ${\rm Cay}(F,B)$ as above, we induce the natural action of $F$ on ${\rm Cay}(F,B)$ to $T_{*}$. This clearly yields also an action of $F$, by automorphisms, on $T$. Since the core $C({\mathcal{V}}_{\mathcal{Q}})$ coincides with the boundary $\partial T$, we equip it with the induced action of $F$ on $\partial T$. Observe that, being label preserving, the above described action of $F$ on $T$ preserves also the lines of the $V$-tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$. As a consequence, the induced action of $F$ on the core ${\mathcal{C}}({\mathcal{V}}_{\mathcal{Q}})=\partial T$ preserves the peripherals in this core.
@@
Let $L$ be a line of the $V$-tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$, and suppose that it corresponds to the generator $w_{i}$ of some subgroup $H_{i}$. Let $P=\partial\gamma_{L}$ be the corresponding peripheral of the core $C({\mathcal{V}}_{\mathcal{Q}})$. Observe that the geodesic $\gamma_{L}$ coincides then, via the identification of $T_{*}$ with ${\rm Cay}(F,B)$, with the axis of the automorphism of ${\rm Cay}(F,B)$ given by left multiplication through some conjugate $gw_{i}g^{-1}$. Since this axis stays at finite Hausdorff distance in ${\rm Cay}(F,B)$ from the subgroup $gH_{i}g^{-1}<F$, it has the same boundary at infinity as this subgroup, and thus we get
@@
Observe that the choices of generators $w_{i}$ in the subgroups $H_{i}<F$ determine orders at all peripherals in the space of peripherals $\partial F$ associated to $\mathcal{Q}$. On the other hand, each $w_{i}$ induces orientations at all edges of any line $L$ in $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$ corresponding to $w_{i}$. It follows that the choices of $w_{i}$ induce orientations at all geodesics $\gamma_{L}$, and thus they induce orders at all peripherals in the core $C({\mathcal{V}}_{\mathcal{Q}})$. It is nor hard to note that the homeomorphism $h$ from the above proof of Lemma 8.19 preserves the above mentioned orders of the peripherals.
@@
Let $F$ be a free finite index subgroup of $\hat{F}$. For each $i\in\{1,\dots,k\}$ and any $\hat{g}\in\hat{F}$ consider the subgroup $\hat{H}_{i}^{\hat{g}}\cap F<F$, which is torsion-free and of finite index in $\hat{H}_{i}^{\hat{g}}$, and which is hence infinite cyclic. Since we clearly have $\hbox{Stab}_{F}\partial\hat{H}_{i}^{\hat{g}}=\hat{H}_{i}^{\hat{g}}\cap F$, this subgroup is a maximal 2-ended subgroup of $F$. Note that $F$ acts on the set $\hat{H}_{i}^{\hat{g}}\cap F:\hat{g}\in\hat{F}/\hat{H}_{i}$ of such subgroups by conjugations, so that
@@
Since $F$ has finite index in $\hat{F}$, the above action by conjugations has finitely many orbits. Denote by $g_{i,j}:1\leq j\leq k_{i}$ some elements of $F$ such that
@@
is a set of representatives of all orbits of the above action of $F$ by conjugations. Since the stabilizer of each such representative satisfies
@@
Then $(F,{\mathcal{P}})$ is obviously an abstract free factor. Since we clearly have $\partial F=\partial\hat{F}$ and $\partial(\hat{H}_{i}^{g_{i,j}})\cap F=\partial\hat{H}_{i}^{g_{i,j}}$, we get the following.
@@
Similarly as in the case of a free abstract factor, the Whitehead graph $W_{\hat{\mathcal{Q}}}$ for $\hat{\mathcal{Q}}$ is not uniquely determined, and depends on the choices of $F$ and $B$. In our later applications of this concept, we will deal with any choice of a subgroup $F$, and such a choice of a basis $B$ for $F$ that the corresponding Whitehaed graph for $(F,{\mathcal{P}})$ is minimal. As we will show, under such choices the Whitehead graphs appearing in our context will be always 2-connected, and in particular essential. Since all further statements concerning the graph $W_{\hat{\mathcal{Q}}}$ apply equally well to all choices as above, we will speak of the Whitehead graph associated to the factor $\hat{\mathcal{Q}}$, meaning any of the graphs as described above.
@@
Next observations concern the $V$-tree system $\Theta[{\mathcal{V}}_{\hat{\mathcal{Q}}}]$ associated to a virtually free abstract factor $\hat{\mathcal{Q}}$.
@@
    Since the subgroups $\hat{H}_{i}^{g_{i,j}}\cap F$ are maximal cyclic in $F$, no two lines in $\Theta[{\mathcal{V}}_{\hat{\mathcal{Q}}}]$ share the set of ends, i.e. for any peripheral $P$ of the core $C({\mathcal{V}}_{\hat{\mathcal{Q}}})$ the set ${\mathcal{L}}_{P}=\{L:\partial\gamma_{L}=P\}$ consists of a single line.
@@
    Since, due to Lemma 8.18, the line $L$ above is orientable in $\Theta[{\mathcal{V}}_{\hat{\mathcal{Q}}}]$, any fixed order of the peripheral $P$ induces orientations at all edges of $W_{\hat{\mathcal{Q}}}$ associated to $P$ (i.e. all edges having appearances in $L$). In accordance with Remark/Definition 7.16(1), we call such orientations order-induced (with respect to the fixed order of $P$).
9. The extended Whitehead graph of a virtually free rigid cluster factor of $G$
Let $v$ be a rigid cluster vertex in the tree $T^{r}_{G}$ of the reduced Bowditch JSJ splitting of a hyperbolic group $G$. We then call the group $G(v)$ a rigid cluster factor of $G$. Denote by $N_{v}$ the set of vertices in $T^{r}_{G}$ adjacent to $v$, and denote by ${\mathcal{H}}_{v}$ the set of the conjugacy classes of the subgroups $G(u)\cap G(v):u\in N_{v}$. Since the family ${\mathcal{H}}_{v}$ is finite, the pair $(G(v),{\mathcal{H}}_{v})$ is an abstract factor, which we denote shortly by ${\mathcal{Q}}_{v}$. Denote also by $(X_{v},{\mathcal{D}}_{v})$ the space with peripherals associated to ${\mathcal{Q}}_{v}$, where ${\mathcal{D}}_{v}$ is the family of peripheral (doubleton) subsets of $X_{v}=\partial G(v)$.
Next lemma exhibits some consequence of Fact 3.4 (which says that the rigid cluster vertices of ${T}^{r}_{G}$ are isolated) for the structure of the reduced Bowditch JSJ splitting of $G$.
@@ 9. The extended Whitehead graph of a virtually free rigid cluster factor of $G$
Suppose on the contrary that $G(u)\cap G(v)$ is a proper subgroup of $G(u)$. This implies that the orbit of the vertex $v$ under the action of $G(u)$ on $T^{r}_{G}$ contains at least one vertex $v^{\prime}\neq v$. By $G$-invariance, this vertex $v^{\prime}$ is also a rigid cluster vertex, and since it is adjacent to $u$, we get a contradiction with the fact that $v$ is isolated (i.e. with Fact 3.4). ∎
@@ 9.1. The Whitehead graph $W_{v}$ of a virtually free rigid cluster factor $G(v)$
Suppose now that the group $G(v)$ (at a rigid cluster vertex $v$ of $T^{r}_{G}$) is a virtually free group, and recall that it is then necessarily non-elementary (see Corollary 3.5). Obviously, ${\mathcal{Q}}_{v}=(G(v),{\mathcal{H}}_{v})$ is then an abstract virtually free factor. Denote by $W_{v}$ any minimal Whitehead graph for ${\mathcal{Q}}_{v}$, as described in Section 8.1 (see Definition 8.23), for any choice of a finite index free subgroup of $G(v)$.
@@
    $W_{\rm min}$ is a 2-connected graph (i.e. connected, and without a cut vertex);
@@
    $F$ has no nontrivial splitting as a free product relative to $\mathcal{H}$.
@@
We say that a vertex $x$ of a graph $X$ is a cut vertex if the space $|X|\setminus\{x\}$ has more connected components than $|X|$. It is an easy observation that any minimal Whitehead graph of a free abstract factor has no cut vertex (see e.g. Lemma 2.3 in [5]). On the other hand, by Proposition 2.6 from [4], the following three conditions are equivalent:
@@
    $F$ splits freely relative to $\mathcal{H}$,
@@
The proposition follows directly from these two observations. ∎
@@
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and let ${\mathcal{Q}}_{v}=(G(v),{\mathcal{H}}_{v})$ be any rigid cluster factor of the Bowditch JSJ splitting for $G$. If the group $G(v)$ is vitrtually free, then any minimal Whitehead graph for ${\mathcal{Q}}_{v}$ (for any choice of a finite index free subgroup $F<G(v)$) is 2-connected.
@@
Suppose on the contrary that some minimal Whitehead graph $W$ for ${\mathcal{Q}}_{v}$ (related to a finite index free subgroup $F$ of $G(v)$, and to the induced from ${\mathcal{H}}_{v}$ family $\mathcal{P}$ of peripheral subgroups of $F$, as described in Section 8.1, right before Lemma 8.22) is not 2-connected. It follows from Proposition 9.2 that $F$ splits freely relative to $\mathcal{P}$. Our next goal is to deduce from this fact the following.
@@
$G(v)$ splits essentially relative to ${\mathcal{H}}_{v}$ over a finite subgroup.
To prove the claim, note that since $G(v)$ is hyperbolic and the subgroups from ${\mathcal{H}}_{v}$ are maximal 2-ended and pairwise nonconjugate, $G(v)$ is relatively hyperbolic with respect to ${\mathcal{H}}_{v}$, and the same is true for the pair $(F,{\mathcal{P}})$. We will argue by referring to some results from the theory of relatively hyperbolic groups. In particular, we will refer to the properties of the so called Bowditch boundary $\partial_{B}(G^{\star},{\mathcal{Q}}^{\star})$ of a relatively hyperbolic group $G^{\star}$ (with respect to a family ${\mathcal{Q}}^{\star}=\{Q_{i}\}$ of subgroups). By result of Tran [25], if $G^{\star}$ is hyperbolic then $\partial_{B}(G^{\star},{\mathcal{Q}}^{\star})$ is homeomorphic to the quotient of the Gromov boundary $\partial G^{\star}$ through a decomposition provided by the family of subspaces
@@
By Proposition 10.1 in [2], the Bowditch boundary $\partial_{B}(G^{\star},{\mathcal{Q}}^{\star})$ is connected iff $G^{\star}$ does not split over a finite subgroup relative to ${\mathcal{Q}}^{\star}$. It follows that the Bowditch boundary $\partial_{B}(F,{\mathcal{P}})$ is not connected. As a consequence, the Bowditch boundary $\partial_{B}(G(v),{\mathcal{H}}_{v})$ (which is homeomorphic to $\partial_{B}(F,{\mathcal{P}})$) is also not connected. By the same Proposition 10.1 in [2], $G(v)$ splits over a finite subgroup relative to ${\mathcal{H}}_{v}$, which completes the proof of the claim.
Coming back to the proof of Proposition 9.3, it is not hard to observe that if $G$ contains a rigid cluster factor which splits over a finite subgroup relative to its peripheral subgroups, the group $G$ itself also splits nontrivially over a finite subgroup (as an amalgameted free product or an HNN-extension). Due to a well known result of Stalllings, the latter implies that $G$ is not 1-ended, contradicting one of our assumptions, and thus the proposition follows. ∎

### v2

@@ 8.3. Abstract virtually free factors and their Whitehead graphs
Recall that the tree $T$ is bipartite, with all white vertices of degree 2. Thus $T$ is the first barycentric subdivision of a certain tree, which we denote $T_{*}$. The vertices of $T_{*}$ are naturally identified with the black vertices of $T$. Our goal is to identify $T_{*}$ with ${\rm Cay}(F,B)$. To do this, we associate to each edge $e$ of $T_{*}$ some orientation and some label from the generating set $B$. Let $t,s$ be the endpoints of $e$, and view them as black vertices of $T$. Let $u$ be the barycenter of $e$, which we view as an appropriate white vertex of $T$ (the unique white vertex of $T$ adjacent to both $t$ and $s$). The constituent spaces at $t$ and $s$ of the tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$, denoted $\Gamma_{t}^{\circ}$ and $\Gamma_{s}^{\circ}$, are the copies of the punctured Whitehead graph $W_{\mathcal{Q}}$. The images in $\Gamma_{t}^{\circ}$ and in $\Gamma_{s}^{\circ}$ of the peripheral space $\Sigma_{u}$ are some $V$-peripherals related to vertices of $\Gamma_{t}$ and $\Gamma_{s}$, and we denote these vertices by $z$ and $y$, respectively. Identifying both $z$ and $y$ as vertices of the Whitehead graph $W_{\mathcal{Q}}$, we have $z,y\in B\cup B^{-1}$, and by the form of the associated $V$-involution $a_{\mathcal{Q}}$, we have $y=z^{-1}$. Without loss of generality, assume that $z\in B$. Orient then the edge $e$ consistently with the order $(z,y)$, and label it with $z$. It is then easy to observe that for any vertex $v$ of $T_{*}$ we have exactly $m$ oriented edges issuing from $v$, labelled bijectively by the elements from $B$, and exactly $m$ oriented edges incoming to $v$, also labelled bijectively with the elements of $B$. Thus the tree $T_{*}$ is locally isomorphic to the Cayley graph ${\rm Cay}(F,B)$, which is also a tree. To get an explicit identification of $T_{*}$ with ${\rm Cay}(F,B)$, fix a vertex of $T_{*}$ and identify it with the vertex of ${\rm Cay}(F,B)$ corresponding to the unit of $F$. This identification obviously extends to a unique label preserving isomorphism of oriented graphs, and we take this isomorphism as the desired identification.
Given the identification of $T_{*}$ with ${\rm Cay}(F,B)$ as above, the natural action of $F$ on ${\rm Cay}(F,B)$ induces an action of $F$ on $T_{*}$. This clearly yields also an action of $F$, by automorphisms, on $T$. Since the core $C({\mathcal{V}}_{\mathcal{Q}})$ coincides with the boundary $\partial T$, we equip it with the induced action of $F$ on $\partial T$. Observe that, being orientation and label preserving, the above described action of $F$ on $T$ preserves also the lines of the $V$-tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$, as explained in some detail in the proof of Lemma 8.19. As a consequence, the induced action of $F$ on the core ${\mathcal{C}}({\mathcal{V}}_{\mathcal{Q}})=\partial T$ preserves the peripherals in this core.
@@
Let $L$ be a line of the $V$-tree system $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$. Observe that any two consecutive edges of $L$ correspond then to some consecutive occurences of subwords of form $xy$ and $yz$ in precisely one of the cyclically reduced words $w_{i}$. We will emphasize this fact by saying that the line $L$ corresponds to this generator $w_{i}$ (of the corresponding peripheral subgroups $H_{i}$). Let $P=\partial\gamma_{L}$ be the corresponding peripheral of the core $C({\mathcal{V}}_{\mathcal{Q}})$. Observe that the geodesic $\gamma_{L}$ coincides then, via the identification of $T_{*}$ with ${\rm Cay}(F,B)$, with the axis of the automorphism of ${\rm Cay}(F,B)$ given by left multiplication through some conjugate $gw_{i}g^{-1}$. Since this axis stays at finite Hausdorff distance in ${\rm Cay}(F,B)$ from the subgroup $gH_{i}g^{-1}<F$, it has the same boundary at infinity as this subgroup, and thus we get
@@
Observe that the choices of the generators $h_{i}$ in the subgroups $H_{i}<F$ determine (uniquely) the choices of the generators $gh_{i}g^{-1}$ in the conjugate subgroups $gH_{i}g^{-1}<F$, and hence also the orders on all peripherals in the space of peripherals $\partial F$ associated to $\mathcal{Q}$. On the other hand, each $w_{i}$ induces orientations on all edges of any line $L$ in $\Theta[{\mathcal{V}}_{\mathcal{Q}}]$ corresponding to $w_{i}$. It follows that the choices of $w_{i}$ induce orientations on all geodesics $\gamma_{L}$, and thus they induce orders on all peripherals in the core $C({\mathcal{V}}_{\mathcal{Q}})$. It is not hard to see that the homeomorphism $h$ from the above proof of Lemma 8.19 preserves the above mentioned orders of the peripherals.
@@
Let $F$ be a free finite index subgroup of $\hat{F}$. For each $i\in\{1,\dots,k\}$ and any $\hat{g}\in\hat{F}$ consider the subgroup $\hat{H}_{i}^{\hat{g}}\cap F<F$, which is torsion-free and of finite index in $\hat{H}_{i}^{\hat{g}}$, and which is hence infinite cyclic. Since we clearly have $\hbox{Stab}_{F}\partial\hat{H}_{i}^{\hat{g}}=\hat{H}_{i}^{\hat{g}}\cap F$, this subgroup is a maximal 2-ended subgroup of $F$. Note that $F$ acts on the set $\{\hat{H}_{i}^{\hat{g}}\cap F:\hat{g}\in\hat{F}/\hat{H}_{i}\}$ of such subgroups by conjugations, as in
@@
Since $F$ has finite index in $\hat{F}$, the above action by conjugations has finitely many orbits. Denote by $g_{i,j}:1\leq j\leq k_{i}$ some elements of $\hat{F}$ such that
@@
is a family of representatives of all orbits of the above action of $F$ by conjugation. Since the stabilizer of each such representative satisfies
@@
Then $(F,{\mathcal{P}})$ is obviously an abstract free factor. Since we clearly have $\partial F=\partial\hat{F}$ and $\partial(\hat{H}_{i}^{g_{i,j}}\cap F)=\partial\hat{H}_{i}^{g_{i,j}}$, we get the following.
@@
Similarly as in the case of a free abstract factor, the Whitehead graph $W_{\hat{\mathcal{Q}}}$ for $\hat{\mathcal{Q}}$ is not uniquely determined, and depends on the choices of $F$ and $B$. In our later applications of this concept, we will deal with any choice of a subgroup $F$, and such a choice of a basis $B$ for $F$ that the corresponding Whitehead graph for $(F,{\mathcal{P}})$ is minimal. As we will show (see Propositions 9.2 and 9.3), under such choices the Whitehead graphs appearing in our context will always be 2-connected, and in particular essential. Since all further statements concerning the graph $W_{\hat{\mathcal{Q}}}$ apply equally well to all choices as above, we will speak of the Whitehead graph associated to the factor $\hat{\mathcal{Q}}$, meaning any of the graphs as described above.
@@
The following observations concern some further features of the $V$-tree system $\Theta[{\mathcal{V}}_{\hat{\mathcal{Q}}}]$ associated to a virtually free abstract factor $\hat{\mathcal{Q}}$.
@@
    Since the subgroups $\hat{H}_{i}^{g_{i,j}}\cap F$ are maximal cyclic in $F$, no two lines in $\Theta[{\mathcal{V}}_{\hat{\mathcal{Q}}}]$ have the same ends, i.e., for any peripheral $P$ of the core $C({\mathcal{V}}_{\hat{\mathcal{Q}}})$ the set ${\mathcal{L}}_{P}=\{L:\partial\gamma_{L}=P\}$ consists of a single line.
@@
    Since, due to Lemma 8.18, the line $L\in\mathcal{L}_{P}$, as above in (1), is orientable in $\Theta[{\mathcal{V}}_{\hat{\mathcal{Q}}}]$, any fixed order of the peripheral $P$ induces orientations on all the edges of $W_{\hat{\mathcal{Q}}}$ associated to $P$ (i.e. all edges having appearances in $L$). In accordance with Remark/Definition 7.16(1), we call such orientations order-induced (with respect to the fixed order of $P$).
9. The (extended) Whitehead graph of a virtually free rigid cluster factor of $G$
Let $v$ be a rigid cluster vertex in the tree $T^{r}_{G}$ of the reduced Bowditch JSJ splitting of a hyperbolic group $G$ (see Section 3). We then call the group $G(v)$ a rigid cluster factor of $G$. Denote by $N_{v}$ the set of vertices in $T^{r}_{G}$ adjacent to $v$, and denote by ${\mathcal{H}}_{v}$ the set of the $G(u)$-conjugacy classes of the subgroups $G(u)\cap G(v):u\in N_{v}$. Since the family ${\mathcal{H}}_{v}$ is finite, the pair $(G(v),{\mathcal{H}}_{v})$ is an abstract factor, which we denote briefly by ${\mathcal{Q}}_{v}$. Denote also by $(X_{v},{\mathcal{D}}_{v})$ the space with peripherals associated to ${\mathcal{Q}}_{v}$, where ${\mathcal{D}}_{v}$ is the family of peripheral (doubleton) subsets of $X_{v}=\partial G(v)$.
The next lemma exhibits some consequence of Fact 3.5 (which says that the rigid cluster vertices of ${T}^{r}_{G}$ are isolated) for the structure of the reduced Bowditch JSJ splitting of $G$.
@@ 9. The (extended) Whitehead graph of a virtually free rigid cluster factor of $G$
Suppose on the contrary that $G(u)\cap G(v)$ is a proper subgroup of $G(u)$. This implies that the orbit of the vertex $v$ under the action of $G(u)$ on $T^{r}_{G}$ contains at least one vertex $v^{\prime}\neq v$. By $G$-invariance, this vertex $v^{\prime}$ is also a rigid cluster vertex, and since it is adjacent to $u$, we get a contradiction with the fact that $v$ is isolated (i.e. with Fact 3.5). ∎
@@ 9.1. The Whitehead graph $W_{v}$ of a virtually free rigid cluster factor $G(v)$
Suppose now that the group $G(v)$ (at a rigid cluster vertex $v$ of $T^{r}_{G}$) is a virtually free group, and recall that it is then necessarily non-elementary (see Corollary 3.7). Obviously, ${\mathcal{Q}}_{v}=(G(v),{\mathcal{H}}_{v})$ is then an abstract virtually free factor. Denote by $W_{v}$ any minimal Whitehead graph for ${\mathcal{Q}}_{v}$, as described in Section 8.3 (see Definition 8.23 and Remark 8.24), for any choice of a finite index free subgroup of $G(v)$.
@@
    $W_{\rm min}$ is a 2-connected graph (i.e. finite, connected, nonempty, not equal to a single vertex, and cutpoint-free);
@@
    $F$ has no splitting as a free product relative to $\mathcal{H}$ (i.e. a splitting as a free product with nontrivial factors where the classes in $\mathcal{H}$ are all elliptic).
@@
We say that a vertex $x$ of a graph $X$ is a cut vertex if the space $|X|\setminus\{x\}$ has more connected components than $|X|$. It is an easy observation that any minimal Whitehead graph of a free abstract factor has no cut vertex (see e.g. Lemma 2.3 in [Cas17]) so any component is either cutpoint-free or an isolated edge. On the other hand, by Proposition 2.6 from [Cas16], the following three conditions are equivalent:
@@
    $F$ splits as a free product relative to $\mathcal{H}$,
@@
The proposition follows directly from these two observations and the fact that $F$ is non-elementary. ∎
@@
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and let ${\mathcal{Q}}_{v}=(G(v),{\mathcal{H}}_{v})$ be any rigid cluster factor of the Bowditch JSJ splitting for $G$. If the group $G(v)$ is virtually free, then any minimal Whitehead graph for ${\mathcal{Q}}_{v}$ (for any choice of a finite index free subgroup $F<G(v)$) is 2-connected.
@@
Suppose on the contrary that some minimal Whitehead graph $W$ for ${\mathcal{Q}}_{v}$ (related to a finite index free subgroup $F$ of $G(v)$, and to the family $\mathcal{P}$ of peripheral subgroups of $F$ induced from ${\mathcal{H}}_{v}$, as described in Section 8.3, right before Lemma 8.22) is not 2-connected. It follows from Proposition 9.2 that $F$ splits as a free product relative to $\mathcal{P}$. Our next goal is to deduce from this fact the following.
@@
$G(v)$ splits relative to ${\mathcal{H}}_{v}$ over a finite subgroup (i.e. acts on a tree without global fixed points such that edge stabilizers are finite and subgroups from ${\mathcal{H}}_{v}$ are conjugate into vertex stabilizers).
To prove the claim, note that since $G(v)$ is hyperbolic and the subgroups from ${\mathcal{H}}_{v}$ are maximal 2-ended and pairwise nonconjugate, $G(v)$ is relatively hyperbolic with respect to ${\mathcal{H}}_{v}$, and the same is true for the pair $(F,{\mathcal{P}})$. We will argue by referring to some results from the theory of relatively hyperbolic groups. In particular, we will refer to the properties of the so called Bowditch boundary $\partial_{B}(G^{\star},{\mathcal{Q}}^{\star})$ of a relatively hyperbolic group $G^{\star}$ (with respect to a family ${\mathcal{Q}}^{\star}=\{Q_{i}\}$ of subgroups). By a result of Tran [Tra13], if $G^{\star}$ is hyperbolic then $\partial_{B}(G^{\star},{\mathcal{Q}}^{\star})$ is homeomorphic to the quotient of the Gromov boundary $\partial G^{\star}$ through a decomposition provided by the family of subspaces
@@
By Proposition 10.1 in [Bow12], the Bowditch boundary $\partial_{B}(G^{\star},{\mathcal{Q}}^{\star})$ is connected iff $G^{\star}$ does not split over a finite subgroup relative to ${\mathcal{Q}}^{\star}$. Since $F$ splits over the trivial group relative to $\mathcal{P}$, it follows that the Bowditch boundary $\partial_{B}(F,{\mathcal{P}})$ is not connected. As a consequence, the Bowditch boundary $\partial_{B}(G(v),{\mathcal{H}}_{v})$ (which is homeomorphic to $\partial_{B}(F,{\mathcal{P}})$) is also not connected. By the same Proposition 10.1 in [Bow12], $G(v)$ splits over a finite subgroup relative to ${\mathcal{H}}_{v}$, which completes the proof of the claim.
Coming back to the proof of Proposition 9.3, it is not hard to observe that if $G$ contains a rigid cluster factor which splits over a finite subgroup relative to its peripheral subgroups, the group $G$ itself also splits nontrivially over a finite subgroup (as an amalgameted free product or an HNN-extension). Due to a well known result of Stallings, the latter implies that $G$ is not 1-ended, contradicting one of our assumptions, and thus the proposition follows. ∎
As we have already mentioned in Remark 8.24, in our further considerations we will use the convention of choosing as “the Whitehead graph $W_{v}$” any minimal Whitehead graph for $\mathcal{Q}_{v}$. In particular, $W_{v}$ will be always 2-connected.